# An Elliott intertwining approach to classifying actions of $\mathrm{C}^{*}$-tensor categories

### 29 Jan 2024 v2

@@ 3. The generalised cocycle category
where $\iota(x)=\lim_{\lambda}x\boxtimes\eta_{\lambda}$ for some approximate unit $\eta_{\lambda}$ of $B$ and all $x\in F(X)$, and $f$ is the map given by $f(b\boxtimes y)=b\rhd y$ for all $b\in B$ and $y\in G(X)$. Note that $f$ is an $A$-$B$-bimodule isomorphism if we see $G(X)$ as a left $A$-module through $\phi$ (i.e. $\phi(a)\rhd_{B}f(y)=a\rhd_{A}f(y)$ for all $y\in{}_{\phi}B\boxtimes G(X)$ and $a\in A$).
To check that $\iota$ is well-defined, we show that the net $x\boxtimes\eta_{\lambda}$ is Cauchy for all $X\in\mathcal{C}$ and $x\in F(X)$. Precisely, one has that $\langle x\boxtimes(\eta_{\lambda}-\eta_{\mu}),x\boxtimes(\eta_{\lambda}-\eta_{\mu})\rangle_{B}=\langle\eta_{\lambda}-\eta_{\mu},\langle x,x\rangle_{A}\rhd(\eta_{\lambda}-\eta_{\mu})\rangle_{B}=|\phi(\langle x,x\rangle_{A})^{1/2}(\eta_{\lambda}-\eta_{\mu})|^{2}$. This converges to 00 since the image of $\phi$ is contained in $B$ and $\eta_{\lambda}$ is an approximate unit for $B$.
We now check that the family $\{h^{X}\}$ defined above satisfies the required compatibility conditions. Since each of the maps in (3.13) are linear, we get that $h^{X}$ is linear. To see (i), note that $f$, $\mathbbm{v}_{X}$ and $\iota$ are left module maps so $\phi(a)\rhd h^{X}(x)=h^{X}(a\rhd x)$ for all $a\in A$ and $x\in F(X)$. Moreover, as $f$ and $\mathbbm{v}_{X}$ are right $B$-module maps and $\iota$ is a right $A$-module map
@@
Now, let $\Phi:\{\mathbbm{v}_{X}\}\to\{h^{X}\}$ be the map induced by the formula in (3.13). Note that $\iota$ and hence $\Phi$ is independent of the choice of approximate unit. Indeed, let $\eta_{\lambda}$ and $\xi_{\lambda}$ be two approximate units for $B$. We show that the net $x\boxtimes(\eta_{\lambda}-\xi_{\lambda})$ converges to 00 for any $x\in F(X)$. Note that
@@
Note that this alternative picture only holds for cocycle morphisms. In the generality of cocycle representations, the maps $h^{X}$ may not be well-defined. That is because if $(\phi,\mathbbm{v})$ is a cocycle representation, $\phi$ can land in $M(B)\setminus B$. As $\eta_{\lambda}$ is an approximate unit for $B$, $\|\phi(\langle x,x\rangle_{A})^{1/2}(\eta_{\lambda}-\eta_{\mu})\|^{2}$ need not converge to 00.
@@
When $\phi:A\rightarrow B$ is extendible and $(\phi,\{h^{X}\}_{X\in\mathcal{C}})$ is a cocycle morphism, condition (i) also follows for $a,a^{\prime}\in\mathcal{M}(A)$. Precisely, if $a,a^{\prime}\in\mathcal{M}(A)$ then $h^{X}(a\rhd x\lhd a^{\prime})=\phi_{p}(a)\rhd h^{X}(x)\lhd\phi_{p}(a^{\prime})$ for any $X\in\mathcal{C}$ and $x\in F(X)$ where $p\in\mathcal{M}(B)$ is the projection associated to $\phi$. Similarly, for any $u,u^{\prime}\in\mathcal{U}(\mathcal{M}(A))$ it follows that $h^{X}(u\rhd x\lhd u^{\prime})=\phi^{\dagger}(u)\rhd h^{X}(x)\lhd\phi^{\dagger}(u^{\prime})$. Indeed, for an approximate unit $e_{\lambda}$ of $A$, $\phi(e_{\lambda})$ converges to $p$ and
@@
So $\phi^{\dagger}(u)\rhd h^{X}(x)\lhd\phi^{\dagger}(u^{\prime})=\phi_{p}(u)\rhd h^{X}(x)\lhd\phi_{p}(u^{\prime})=h^{X}(u\rhd x\lhd u^{\prime})$.
@@
Suppose $\mathcal{C}=\mathrm{Hilb}(\Gamma)$ and $(\phi,\{h^{g}\}_{g\in\Gamma}):(A,\alpha,J)\to(B,\beta,I)$ is an extendible cocycle morphism. Recall from Example 3.3 that $\mathbbm{v}_{g}(a\boxtimes b)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes\mathbbm{u}_{g}\phi(a)b,$ where $\eta_{\lambda}$ is an approximate unit for $B$ and $(\mathbbm{u}_{g})_{g\in\Gamma}\subseteq\mathcal{M}(B)$. Then for any $g\in\Gamma$ and $a\in A$,
$h^{g}(a)=f\left(\mathbbm{v}_{g}\left(\lim\limits_{\lambda}a\boxtimes\eta_{\lambda}\right)\right)=f\left(\lim\limits_{\lambda}\eta_{\lambda}\boxtimes\mathbbm{u}_{g}\phi(a)\right)=\mathbbm{u}_{g}\phi(a).$
For convenience, we will denote a cocycle morphism by $(\phi,h)$, where $h$ denotes the collection of linear maps $\{h^{X}\}_{X\in\mathcal{C}}$. We now discuss their composition.
@@
Clearly $\psi\circ\phi:A\to C$ is a $*{}^{*}$-homomorphism and $\{l^{X}\circ h^{X}:F(X)\to H(X)\}_{X\in\mathcal{C}}$ is a family of linear maps. Conditions (i), (ii), (iii), and (v) are immediate. The compatibility with the tensor product follows by stacking the diagrams in (iv) of Lemma 3.11 for $h$ and $l$. Thus, $(\psi\circ\phi,l\circ h):(A,F,J)\to(C,H,K)$ induces a cocycle morphism by Lemma 3.11.
@@
where $(\mathbbm{w}\circ\mathbbm{v})_{X}=(\mathrm{id}_{\phi}\boxtimes\mathbbm{w}_{X})\circ(\mathbbm{v}_{X}\boxtimes\mathrm{id}_{\psi})$ (see (3.3)). Start with an elementary tensor $x\boxtimes c$ and let $\zeta_{\mu}$ be an approximate unit of $A$ and $\eta_{\lambda}$ be an approximate unit of $B$. Following the composition of maps defining $(\mathbbm{w}*\mathbbm{v})_{X}$, we get that
@@
Lemma 3.11 shows that any cocycle morphism can be equivalently represented by a pair $(\phi,h)$. Moreover Lemma 3.15 shows that in the latter picture the composition of cocycle morphisms translates to the composition of the underlying $*{}^{*}$-homomorphisms and linear maps. From now on we will freely identify these two pictures.
@@
In the spirit of [40, Definition 1.16], we have constructed a category of $\mathcal{C}$-$\mathrm{C}^{*}$-algebras.
@@ 4. Inductive limits
In this section, we construct inductive limits in $\mathrm{C}^{*}_{\mathcal{C}}$ for semisimple $\mathcal{C}$. This is done in [6, Proposition 4.4], when restricted to unital $\mathrm{C}^{*}$-algebras and unital, injective connecting maps. Our approach is slightly different and does not need these assumptions. Before starting our construction, let us set up some notation.
@@
$A^{(0)}=\left\{(a_{n})_{n\geq 1}\in\frac{\bigoplus_{\ell^{\infty}}A_{n}}{\bigoplus_{c_{0}}A_{n}}:\exists N\in\mathbb{N}\ \text{s.t.}\ \forall n\geq N,\phi_{n}(a_{n})=a_{n+1}\right\}$
@@
for all $n\geq 1$, where we adopt the standard notation $\phi_{n,m}:=\phi_{m-1}\circ\ldots\circ\phi_{n}$ for any $m>n\geq 1$.
@@
To build an action on $A$, we start by constructing bimodules that will form the image of the functor. Essentially, for any $X\in\mathcal{C}$, we can build a Hilbert $A$-$A$-bimodule as an inductive limit of $F_{n}(X)$. The construction is very similar to the one in (2.4).
@@
$F^{(0)}(X)=\Big{\{}(x_{n})_{n\geq 1}\in\frac{\bigoplus_{\ell^{\infty}}F_{n}(X)}{\bigoplus_{c_{0}}F_{n}(X)}:\exists N\in\mathbb{N}\ \text{s.t.}\ \forall n\geq N,h_{n}^{X}(x_{n})=x_{n+1}\Big{\}},$
@@
That (4.4) is independent of the choice of representative sequences, follows as in the proof of Lemma 2.4. Thus, (4.4) gives a well-defined $A^{(0)}$-valued map. It is now straightforward to check that this function is right linear, left conjugate linear, and antisymmetric. Finally, it is clear that $\langle x,x\rangle_{A^{(0)}}\geq 0$ and $\langle x,x\rangle_{A^{(0)}}=0$ if and only if $\langle x_{n},x_{n}\rangle_{A_{n}}$ converges to 00 i.e. $x=0$ in $F^{(0)}(X)$. Thus, the conclusion follows. ∎
For any $X\in\mathcal{C}$, since $A^{(0)}$ is a dense $*{}^{*}$-subalgebra of $A$, combining Lemma 4.1 and [37, Lemma 2.16], we form the completion of $F^{(0)}(X)$, denoted by $F(X)$. This is a right-Hilbert $A$-module.
@@
For any $X\in\mathcal{C}$, with the notation above, $F(X)$ is a non-degenerate right Hilbert $A$-$A$-bimodule.
@@
We start by defining a left $A$-action on $F(X)$. For any $x=(x_{n})_{n\geq 1}\in F^{(0)}(X)$ and any $a=(a_{n})_{n\geq 1}\in A^{(0)}$, we can define
@@
The fact that (4.5) is a well-defined left $A^{(0)}$-action on $F^{(0)}(X)$ follows as in the proof of Lemma 4.1. Moreover, the left $A^{(0)}$-action is adjointable, since the left action of $A_{n}$ on $F_{n}(X)$ is adjointable for any $n\geq 1$. Thus, using [37, Lemma 2.16], it extends to a left action of $A$ on $F(X)$ by density. ∎
We now show that the assignment $X\to F(X)$ extends to an action of $\mathcal{C}$ on $A$.
@@
With the notation above, $F:\mathcal{C}\to\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$ is a $\mathrm{C}^{*}$-functor. Moreover, there exists a unitary natural isomorphism
@@
such that the pair $(F,J)$ is an action of $\mathcal{C}$ on $A$.
@@
Suppose $X,Y\in\mathcal{C}$ and $f:X\to Y$ is a morphism in $\mathcal{C}$. We define $F(f)(x)=(F_{n}(f)(x_{n}))_{n\geq 1}$ for any $x\in F^{(0)}(X)$. As $\mathcal{C}$ is semisimple, by decomposing $X$ and $Y$ into irreducibles one has that $f=\bigoplus\lambda_{i}\mathrm{id}_{X_{i}}$ for finitely many non-zero scalars $\lambda_{i}$ and $X_{i}\in\mathrm{Irr}(\mathcal{C})$. Therefore, as each $F_{n}$ is a $*{}^{*}$-functor, $F_{n}(f)$ is uniformly bounded by the supremum of $\lambda_{i}$ and so $F(f)$ is bounded on $F^{(0)}(X)$. Therefore, one may extend $F(f)$ to a bounded $A$-$A$-bimodule map on $F(X)$. This operator is further adjointable as each $F_{n}(f)$ is adjointable. It is straightforward to check that $F$ defines a $\mathrm{C}^{*}$-functor.
@@
We claim that $J_{X,Y}$ is well-defined. Firstly, note that $J_{X,Y}$ is independent of the choice of representative sequences by continuity of $J_{X,Y}^{(n)}$ for any $n\in\mathbb{N}$.
@@
Moreover, since each $J^{(n)}$ is a unitary natural isomorphism, so is $J$. Finally, since each family $\{J_{X,Y}^{(n)}:X,Y\in\mathcal{C}\}$ satisfies a commuting diagram as in (2.1), so does the collection of maps $\{J_{X,Y}:X,Y\in\mathcal{C}\}$. Hence, the pair $(F,J)$ gives an action of $\mathcal{C}$ on $A$. ∎
@@
For any $n\geq 1$, we have a $*{}^{*}$-homomorphism $\phi_{n,\infty}:A_{n}\to A$. To define a cocycle morphism from $(A_{n},F_{n},J^{(n)})$ to $(A,F,J)$, it suffices to find a collection of linear maps $\{h_{n,\infty}^{X}:F_{n}(X)\to F(X)\}_{X\in\mathcal{C}}$ satisfying the conditions in Lemma 3.11. If $X\in\mathcal{C}$ and $x_{n}\in F_{n}(X)$ for $n\in\mathbb{N}$, we define
@@
We now check that the triple $(A,F,J)$ is the inductive limit of the sequence of triples $(A_{n},F_{n},J^{(n)})$, together with connecting morphisms $(\phi_{n},h_{n})$. Firstly, we claim that (4.6) commutes. Since $A$ is an inductive limit for the system $(A_{n},\phi_{n})_{n\geq 1}$, it follows that for any $n\geq 1$, $\phi_{n,\infty}=\phi_{n+1,\infty}\circ\phi_{n}$. Therefore, it suffices to check that for any $X\in\mathcal{C}$ and any $n\geq 1$, we have that $h_{n,\infty}^{X}=h_{n+1,\infty}^{X}\circ h_{n}^{X}$. This follows by (4.7) since for any $k>n+1$, $h_{n,k}^{X}=h_{n+1,k}^{X}\circ h_{n}^{X}$.
It remains to check that $(A,F,J)$ satisfies the universal property. Precisely, let $B$ be a $\mathrm{C}^{*}$-algebra and $(B,G,I)$ defining an action of $\mathcal{C}$ on $B$. Suppose there exists a sequence of cocycle morphisms $(\psi,l)_{n,\infty}:(A_{n},F_{n},J^{(n)})\to(B,G,I)$ such that the diagram
@@
By the universal property of the inductive limit $A$, there exists a unique $*{}^{*}$-homomorphism $\Phi:A\to B$ such that $\Phi\circ\phi_{n,\infty}=\psi_{n,\infty}$ for all $n\geq 1$. Then, for any $X\in\mathcal{C}$, the union $\bigcup_{k\geq 1}h_{k,\infty}^{X}(F_{k}(X))$ is dense in $F(X)$. Since $l_{k,\infty}^{X}$ is uniformly bounded in $k$, (4.8) yields that $\mathrm{Ker}\ h_{k,\infty}^{X}\subseteq\mathrm{Ker}\ l_{k,\infty}^{X}$ for any $X\in\mathcal{C}$ and any $k\in\mathbb{N}$. Therefore, for any $x_{k}\in F_{k}(X)$, define $r^{X}(h_{k,\infty}^{X}(x_{k}))=l_{k,\infty}^{X}(x_{k})$. As $l_{k,\infty}^{X}$ is uniformly bounded in $k$, we may extend $r^{X}$ to a well-defined linear map $r^{X}:F(X)\to G(X)$. The fact that $r^{X}$ satisfies the conditions appearing in Lemma 3.11 is routine and essentially follows from the fact that for each $k\geq 1$, $(\psi_{k,\infty},l_{k,\infty})$ and $(\phi_{k,\infty},h_{k,\infty})$ are cocycle morphisms. For the convenience of the reader, we check condition (iii) of Lemma 3.11. Let $X\in\mathcal{C}$, $x_{k},y_{k}\in F_{k}(X)$. Then,
$\langle r^{X}(h_{k,\infty}^{X}(x_{k})),r^{X}(h_{k,\infty}^{X}(y_{k}))\rangle$ $=\langle l_{k,\infty}^{X}(x_{k}),l_{k,\infty}^{X}(y_{k})\rangle$
$=\psi_{k,\infty}(\langle x_{k},y_{k}\rangle)$
$=\Phi\circ\phi_{k,\infty}(\langle x_{k},y_{k}\rangle)$
$=\Phi(\langle h_{k,\infty}^{X}(x_{k}),h_{k,\infty}^{X}(y_{k})\rangle).$
The remaining conditions follow similarly. Therefore, $(\Phi,r)$ is the unique cocycle morphism such that $(\Phi,r)\circ(\phi,h)_{n,\infty}=(\psi,l)_{n,\infty}$. Hence, the triple $(A,F,J)$ is the inductive limit of the system in (4.1). ∎
@@
4.1. Topology induced on cocycle representations
[…]

### 08 Apr 2026 v3

@@ 3. The generalised cocycle category
where $\iota(x)=\lim_{\lambda}x\boxtimes\eta_{\lambda}$ for some approximate unit $\eta_{\lambda}$ of $B$ and all $x\in F(X)$, and $f$ is the map given by $f(b\boxtimes y)=b\rhd y$ for all $b\in B$ and $y\in G(X)$. Note that $f$ is an AA-$B$-bimodule isomorphism if we see $G(X)$ as a left AA-module through $\phi$ (i.e. $\phi(a)\rhd_{B}f(y)=a\rhd_{A}f(y)$ for all $y\in{}_{\phi}B\boxtimes G(X)$ and $a\in A$).
To check that $\iota$ is well-defined, we show that the net $x\boxtimes\eta_{\lambda}$ is Cauchy for all $X\in\mathcal{C}$ and $x\in F(X)$. Precisely, one has that $\langle x\boxtimes(\eta_{\lambda}-\eta_{\mu}),x\boxtimes(\eta_{\lambda}-\eta_{\mu})\rangle_{B}=\langle\eta_{\lambda}-\eta_{\mu},\langle x,x\rangle_{A}\rhd(\eta_{\lambda}-\eta_{\mu})\rangle_{B}=|\phi(\langle x,x\rangle_{A})^{1/2}(\eta_{\lambda}-\eta_{\mu})|^{2}$. This converges to 0 since the image of $\phi$ is contained in $B$ and $\eta_{\lambda}$ is an approximate unit for $B$.
We now check that the family $\{h^{X}\}$ defined above satisfies the required compatibility conditions. Since each of the maps in (3.8) are linear, we get that $h^{X}$ is linear. To see (i), note that $f$, $\mathbbm{v}_{X}$ and $\iota$ are left module maps so $\phi(a)\rhd h^{X}(x)=h^{X}(a\rhd x)$ for all $a\in A$ and $x\in F(X)$. Moreover, as $f$ and $\mathbbm{v}_{X}$ are right $B$-module maps and $\iota$ is a right AA-module map
@@
Now, let $\Phi:\{\mathbbm{v}_{X}\}\to\{h^{X}\}$ be the map induced by the formula in (3.8). Note that $\iota$ and hence $\Phi$ is independent of the choice of approximate unit. Indeed, let $\eta_{\lambda}$ and $\xi_{\lambda}$ be two approximate units for $B$. We show that the net $x\boxtimes(\eta_{\lambda}-\xi_{\lambda})$ converges to 0 for any $x\in F(X)$. Note that
@@
Note that this alternative picture only holds for cocycle morphisms. In the generality of cocycle representations, the maps $h^{X}$ may not be well-defined. That is because if $(\phi,\mathbbm{v})$ is a cocycle representation, $\phi$ can land in $M(B)\setminus B$. As $\eta_{\lambda}$ is an approximate unit for $B$, $\|\phi(\langle x,x\rangle_{A})^{1/2}(\eta_{\lambda}-\eta_{\mu})\|^{2}$ need not converge to 0.
Remark 3.11.
Note that condition (i) follows from (iv) as $J_{1,X}$ and $J_{X,1}$ correspond to the left and right actions of AA on $F(X)$. Similarly from (iv) it is clear that $h^{1_{\mathcal{C}}}$ is a -homomorphism so we may simply define a cocycle morphism satisfying conditions (ii)-(iv) by setting $h^{1_{\mathcal{C}}}=\phi$.
@@
Lemma 3.10 shows that any cocycle morphism can be equivalently represented by a pair $(\phi,h)$, where for convenience, we denote a cocycle morphism by $(\phi,h)$, where $h$ denotes the collection of linear maps $\{h^{X}\}_{X\in\mathcal{C}}$. Moreover, Lemma 3.15 below shows that in the latter picture, the composition of cocycle morphisms translates to the composition of the underlying -homomorphisms and linear maps. From now on, we will freely identify these two pictures.
@@
When $\phi:A\rightarrow B$ is extendible and $(\phi,\{h^{X}\}_{X\in\mathcal{C}})$ is a cocycle morphism, condition (i) also follows for $a,a^{\prime}\in\mathcal{M}(A)$. Precisely, if $a,a^{\prime}\in\mathcal{M}(A)$ then $h^{X}(a\rhd x\lhd a^{\prime})=\phi_{p}(a)\rhd h^{X}(x)\lhd\phi_{p}(a^{\prime})$ for any $X\in\mathcal{C}$ and $x\in F(X)$ where $p\in\mathcal{M}(B)$ is the projection associated to $\phi$. Similarly, for any $u,u^{\prime}\in\mathcal{U}(\mathcal{M}(A))$ it follows that $h^{X}(u\rhd x\lhd u^{\prime})=\phi^{\dagger}(u)\rhd h^{X}(x)\lhd\phi^{\dagger}(u^{\prime})$. Indeed, for an approximate unit $e_{\lambda}$ of AA, $\phi(e_{\lambda})$ converges to $p$ and
@@
Likewise, we have $h^{X}(x)\lhd p=h^{X}(x)$. Therefore,
$\phi^{\dagger}(u)\rhd h^{X}(x)\lhd\phi^{\dagger}(u^{\prime})=\phi_{p}(u)\rhd h^{X}(x)\lhd\phi_{p}(u^{\prime})=h^{X}(u\rhd x\lhd u^{\prime}).$
@@
Suppose $(A,\alpha)$ and $(B,\beta)$ are actions of a countable discrete group $\Gamma$ on $\mathrm{C}^{*}$-algebras AA and $B$. Consider them as actions of $\mathrm{Hilb}(\Gamma)$ as in Example 2.4 and let $(\phi,\{h^{g}\}_{g\in\Gamma}):(A,\alpha)\to(B,\beta)$ be an extendible cocycle morphism. Recall from Example 3.3 that $\mathbbm{v}_{\mathbb{C}g}(a\boxtimes b)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes\mathbbm{u}_{g}\phi(a)b,$ where $\eta_{\lambda}$ is an approximate unit for $B$ and $(\mathbbm{u}_{g})_{g\in\Gamma}\subseteq\mathcal{M}(B)$. Then for any $g\in\Gamma$ and $a\in A$,
$h^{g}(a)=f\left(\mathbbm{v}_{\mathbb{C}g}\left(\lim\limits_{\lambda}a\boxtimes\eta_{\lambda}\right)\right)=f\left(\lim\limits_{\lambda}\eta_{\lambda}\boxtimes\mathbbm{u}_{g}\phi(a)\right)=\mathbbm{u}_{g}\phi(a).$
We now discuss the composition of cocycle morphisms.
@@
Clearly $\psi\circ\phi:A\to C$ is a -homomorphism and $\{l^{X}\circ h^{X}:F(X)\to H(X)\}_{X\in\mathcal{C}}$ is a family of linear maps. Conditions (i), (ii), (iii), and (v) are immediate. The compatibility with the tensor product follows by stacking the diagrams in (iv) of Lemma 3.10 for $h$ and $l$. Thus, $(\psi\circ\phi,l\circ h):(A,F,J)\to(C,H,K)$ induces a cocycle morphism by Lemma 3.10.
@@
where $(\mathbbm{w}\circ\mathbbm{v})_{X}=(\mathrm{id}_{\phi}\boxtimes\mathbbm{w}_{X})\circ(\mathbbm{v}_{X}\boxtimes\mathrm{id}_{\psi})$ (see (3.3)). Start with an elementary tensor $x\boxtimes c$ and let $\zeta_{\mu}$ be an approximate unit of AA and $\eta_{\lambda}$ be an approximate unit of $B$. Following the composition of maps defining $(\mathbbm{w}*\mathbbm{v})_{X}$, we get that
@@
In the spirit of [41, Definition 1.16], we have constructed a category of $\mathcal{C}$-$\mathrm{C}^{*}$-algebras.
@@ 4. Inductive limits
In this section, we construct inductive limits in $\mathrm{C}^{*}_{\mathcal{C}}$ for semisimple $\mathcal{C}$. This is done in [7, Proposition 4.4], when restricted to unital $\mathrm{C}^{*}$-algebras and unital, injective connecting maps. Our approach is slightly different and does not need these assumptions. Before starting our construction, let us set up some notation.
@@
$A^{(0)}=\frac{\left\{(a_{n})_{n\geq 1}\in\bigoplus_{\ell^{\infty}}A_{n}:\lim_{n\to\infty}\|\phi_{n}(a_{n})-a_{n+1}\|=0\right\}}{\bigoplus_{c_{0}}A_{n}}$
@@
for all $n\geq 1$, where we adopt the standard notation $\phi_{n,m}:=\phi_{m-1}\circ\ldots\circ\phi_{n}$ and $\phi_{n,n}=\mathrm{id}_{A_{n}}$ for any $m>n\geq 1$.
@@
To build an action on AA, we start by constructing bimodules that will form the image of the functor. Essentially, for any $X\in\mathcal{C}$, we can build a Hilbert AA-AA-bimodule as an inductive limit of $F_{n}(X)$. The construction is very similar to the one in (2.5).
@@
$F^{(0)}(X)=\frac{\{(x_{n})_{n\geq 1}\in\bigoplus_{\ell^{\infty}}F_{n}(X):\lim_{n\to\infty}\|h_{n}^{X}(x_{n})-x_{n+1}\|=0\}}{\bigoplus_{c_{0}}F_{n}(X)},$
@@
That (4.4) is independent of the choice of representative sequences, follows as in the proof of Lemma 2.6. Thus, (4.4) gives a well-defined $A^{(0)}$-valued map. It is now straightforward to check that this function is right linear, left conjugate linear, and antisymmetric. Finally, it is clear that $\langle x,x\rangle_{A^{(0)}}\geq 0$ and $\langle x,x\rangle_{A^{(0)}}=0$ if and only if $\langle x_{n},x_{n}\rangle_{A_{n}}$ converges to 0 i.e. $x=0$ in $F^{(0)}(X)$. Thus, the conclusion follows. ∎
For any $X\in\mathcal{C}$, since $A^{(0)}$ is a dense -subalgebra of AA, combining Lemma 4.1 and [39, Lemma 2.16], we form the completion of $F^{(0)}(X)$, denoted by $F(X)$. This is a right-Hilbert AA-module.
@@
For any $X\in\mathcal{C}$, with the notation above, $F(X)$ is a non-degenerate right Hilbert AA-AA-bimodule.
@@
We start by defining a left AA-action on $F(X)$. For any $x=(x_{n})_{n\geq 1}\in F^{(0)}(X)$ and any $a=(a_{n})_{n\geq 1}\in A^{(0)}$, we can define
@@
The fact that (4.5) is a well-defined left $A^{(0)}$-action on $F^{(0)}(X)$ follows as in the proof of Lemma 4.1. Moreover, the left $A^{(0)}$-action is adjointable, since the left action of $A_{n}$ on $F_{n}(X)$ is adjointable for any $n\geq 1$. Thus, using [39, Lemma 2.16], it extends to a left action of AA on $F(X)$ by density. ∎
We now show that the assignment $X\to F(X)$ extends to an action of $\mathcal{C}$ on AA.
@@
With the notation above, $F:\mathcal{C}^{\mathrm{rev}}\to\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$ is a $\mathrm{C}^{*}$-functor. Moreover, there exists a unitary natural isomorphism
@@
such that the pair $(F,J)$ is an action of $\mathcal{C}$ on AA.
@@
Suppose $X,Y\in\mathcal{C}$ and $f:X\to Y$ is a morphism in $\mathcal{C}$. We define $F(f)(x)=(F_{n}(f)(x_{n}))_{n\geq 1}$ for any $x\in F^{(0)}(X)$. As each $F_{n}$ is a C-functor, the assignment $f\mapsto F_{n}(f)$ is contractive (see for example [20, Definition 1.5]) and so $F(f)$ is bounded on $F^{(0)}(X)$. Therefore, one may extend $F(f)$ to a bounded AA-AA-bimodule map on $F(X)$. This operator is further adjointable as each $F_{n}(f)$ is adjointable. It is straightforward to check that $F$ defines a $\mathrm{C}^{*}$-functor.
@@
We claim that $J_{X,Y}$ is well-defined. Firstly, note that $J_{X,Y}$ is independent of the choice of representative sequences as $J_{X,Y}^{(n)}$ is an isometry for any $n\in\mathbb{N}$ and $\lim_{n\to\infty}(F_{n}(Y)\boxtimes F_{n}(X))\cong F(Y)\boxtimes F(X)$.
@@
Moreover, since each $J^{(n)}$ is a unitary natural isomorphism, so is $J$. Finally, since each family $\{J_{X,Y}^{(n)}:X,Y\in\mathcal{C}\}$ satisfies a commuting diagram as in (2.1), so does the collection of maps $\{J_{X,Y}:X,Y\in\mathcal{C}\}$. Hence, the pair $(F,J)$ gives an action of $\mathcal{C}$ on AA. ∎
@@
Recalling (4.2), for any $n\geq 1$, we have a -homomorphism $\phi_{n,\infty}:A_{n}\to A$. To define a cocycle morphism from $(A_{n},F_{n},J^{(n)})$ to $(A,F,J)$, it suffices to find a collection of linear maps $\{h_{n,\infty}^{X}:F_{n}(X)\to F(X)\}_{X\in\mathcal{C}}$ satisfying the conditions in Lemma 3.10. If $X\in\mathcal{C}$ and $x_{n}\in F_{n}(X)$ for $n\in\mathbb{N}$, we define
@@
We now check that the triple $(A,F,J)$ is the inductive limit of the sequence of triples $(A_{n},F_{n},J^{(n)})$, together with connecting morphisms $(\phi_{n},h_{n})$. Firstly, it is clear that (4.6) commutes. It remains to check that $(A,F,J)$ satisfies the universal property. Precisely, let $B$ be a $\mathrm{C}^{*}$-algebra and $(B,G,I)$ defining an action of $\mathcal{C}$ on $B$. Suppose there exists a sequence of cocycle morphisms $(\psi,l)_{n,\infty}:(A_{n},F_{n},J^{(n)})\to(B,G,I)$ such that the diagram
@@
For any $X\in\mathcal{C}$, the union $\bigcup_{k\geq 1}h_{k,\infty}^{X}(F_{k}(X))$ is dense in $F(X)$. Since $l_{k,\infty}^{X}$ is contractive, (4.8) yields that $\mathrm{Ker}\ h_{k,\infty}^{X}\subseteq\mathrm{Ker}\ l_{k,\infty}^{X}$ for any $X\in\mathcal{C}$ and any $k\in\mathbb{N}$. Therefore, for any $x_{k}\in F_{k}(X)$, define $r^{X}(h_{k,\infty}^{X}(x_{k}))=l_{k,\infty}^{X}(x_{k})$. As $l_{k,\infty}^{X}$ is contractive, we may extend $r^{X}$ to a well-defined linear map $r^{X}:F(X)\to G(X)$. The fact that $r^{X}$ satisfies the conditions appearing in Lemma 3.10 is routine and essentially follows from the fact that for each $k\geq 1$, $(\psi_{k,\infty},l_{k,\infty})$ and $(\phi_{k,\infty},h_{k,\infty})$ are cocycle morphisms. Therefore, $(\Phi,r)$ is the unique cocycle morphism such that $(\Phi,r)\circ(\phi,h)_{n,\infty}=(\psi,l)_{n,\infty}$. Hence, the triple $(A,F,J)$ is the inductive limit of the system in (4.1). ∎
@@
5. Approximate unitary equivalence
Let $\mathcal{C}$ be a semisimple $\mathrm{C}^{*}$-tensor category with countably many isomorphism classes of simple objects. To define approximate unitary equivalence, we first introduce a topology on the space of cocycle representations between two actions $\mathcal{C}\overset{F}{\curvearrowright}A$ and $\mathcal{C}\overset{G}{\curvearrowright}B$. For this we employ the same approach as in [41, Section 2.1]; introducing a family of pseudometrics which measure the distance between cocycle representations.